Please help with fractions using the distributive property on both sides of the equation
1/4(3x+4)=1/3(x-12)

Answer by stanbon(75887)

1/4(3x+4)=1/3(x-12)
-------
(3/4)x +1 = (1/3)x - 4
---------
[3/4 - (1/3)x = -5
---------
[(9-4)/12]x = -5
---------
(5/12)x = -5
---------
x = (12/5)(-5)
-----
x = 12*-1
-----
x = -12
==============
Cheers,
Stan H.